OpenStudy (anonymous):

X + 4 < - 9 x < - 13 how do you solve it?

OpenStudy (shadowfiend):

Did you mean to make the X uppercase and the x lowercase?

OpenStudy (shadowfiend):

And if so, are you solving for X or x?

OpenStudy (anonymous):

x

OpenStudy (shadowfiend):

Ok. In that case, divide everything by -9: \begin{align} X + 4 &< - 9 x &< - 13\\\\ \frac{X + 4}{-9} &< \frac{-9x}{-9} &< \frac{-13}{-9}\\ \end{align} Because you're dividing by a negative, you need to flip the inequality, so the <s become >s: \[\frac{X+4}{-9} > x > \frac{13}{9}\]